Company ABC has been trading within the UK for the past 5 years. In its first year of operations, it suffered a loss of $5,000. Post that, it has made profits of $6,000 in Year 2, $10,000 in Year 3, $15,000 in Year 4 and $20,000 in Year 5. Calculate the corporate tax in each year of operation for Company ABC, if the corporation tax rate is fixed at 19%
Answer #1
Year Profit corporate tax
1 - 5,000 - 5,000 x 19% = $950 (Tax savings)
2 6,000 6,000 x 19% = 1,140 - 950 = 190
3 10,000 10,000 x 19% = 1,900
4 15,000 15,000 x 19% = 2,850
5 20,000 20,000 x 19% = 3,800

Please ask if you have any query related to the question. Thank you